OpenStudy (kainui):

What is the cube root of -1?

OpenStudy (anonymous):

-1.

OpenStudy (anonymous):

Taking an odd root of a number will always give you an answer. Taking an even root of a number will give you an answer only for a positive number. Negative numbers wouldn't work unless you're doing complex numbers.

OpenStudy (kainui):

Not entirely what I'm looking for, you see that link there? The real answer is:\[\frac{ 1 }{ 2}+i \frac{ \sqrt{3} }{ 2}\] So what's the major distinction here that wolfram alpha would say that cube root of -1 is not -1?

OpenStudy (anonymous):

I didn't get that. It said -1.

OpenStudy (kainui):

I typed that in, and it said, "False".

hartnn (hartnn):

there are 3 cube roots of -1 -1 and 2 complex roots. roots of the equation x^3+1=0

OpenStudy (anonymous):

Ah okay. And yeah hartnn beat me to it.

hartnn (hartnn):

wolf goes insane sometimes and produces incorrect/complicated answers...

OpenStudy (kainui):

Fair enough. It's just a computer, just making sure I wasn't missing some interesting detail of mathematics going on behind the scenes.